ZombieSlayer:

What is the inverse of the function f(x) = one-ninthx + 2? h(x) = 18x – 2 h(x) = 9x – 18 h(x) = 9x + 18 h(x) = 18x + 2

TWALKER16:

what do u think

ZombieSlayer:

I think it's B, but I aso second guess muself

TWALKER16:

yes

ZombieSlayer:

oh, thank you!

TWALKER16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieSlayer I think it's B, but I aso second guess muself \(\color{#0cbb34}{\text{End of Quote}}\) it is

TWALKER16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieSlayer oh, thank you! \(\color{#0cbb34}{\text{End of Quote}}\) np

justus:

To find the inverse, switch the x and y and then solve for y. x = 1/9 y + 2 x − 2 = 1/9 y y = 9x − 18